# Multifractal critical phase driven by coupling quasiperiodic systems to electromagnetic cavities

### v1

@@ Abstract
We theoretically investigate the effects of criticality and multifractal states in a one-dimensional Aubry-Andre-Harper model coupled to electromagnetic cavities. We focus on two specific cases where the phonon frequencies are $\omega_{0}=1$ and $\omega_{0}=2$, respectively. Phase transitions are analyzed using both the average and minimum inverse participation ratio to identify metallic, fractal, and insulating states. We provide numerical evidence to show that the presence of the optical cavity induces a critical, intermediate phase in between the extended and localized phases, hence drastically modifying the traditional transport phase diagram of the Aubry-Andre-Harper model, in which critical states can only exist at the well-defined metal-insulator critical point. We also investigate the probability distribution of the inverse participation ratio and conduct a multifractal analysis to characterize the nature of the critical phase, in which we show that extended, localized, and fractal eigenstates coexist. Altogether our findings reveal the pivotal role that the coupling to electromagnetic cavities plays in tailoring critical transport phenomena at the microscopic level of the eigenstates.
@@ I Introduction
In view of these stimulating results, a great theoretical effort has been made over the past years, to understand the leading effects of non-perturbative cavity light-matter coupling. For instance, Ref. Hagenmüller et al., 2017 showed that the coupling with the cavity enhances the conductivity, while a theoretical approach for a 2D electron gas with Rashba spin-orbit coupling in a cavity QED is presented in Ref. Nataf et al., 2019. Other important theoretical developments were also made for topological phases Guerci et al. (2020); Jangjan and Hosseini (2020); Dmytruk and Schirò (2022); Allard and Weick (2023); Ezawa (2024); Liu et al. (2023); Nguyen et al. (2024), Majorana fermions Bacciconi et al. (2024); Gómez-León et al. (2024), entanglement properties Passetti et al. (2023), magnetic properties Ashida et al. (2020); Masuki and Ashida (2024), Kondo effect Mochida and Ashida (2024), and disordered systems Arwas and Ciuti (2023); Svintsov et al. (2024). In particular, Ref. Svintsov et al. (2024) investigated the effects of cavity QED in an insulating 1D disordered chain, demonstrating that conductivity can be enhanced by several orders of magnitude due to the emission and absorption of virtual photons. This implies that the localization length is highly dependent on the coupling with the cavity Svintsov et al. (2024). Since the 1D Anderson model does not exhibit a metal-insulator transition, these results only apply to exponentially localized, Anderson wavefunctions in disordered media. However, the impact of the coupling to electromagnetic cavities on metal-insulator transitions remains unknown.
@@
To address this issue, we consider a simple Hamiltonian exhibiting extended-localized phase transition, the Aubry-Andre-Harper (AAH) model Harper (1955); Aubry and André (1980). It describes an one-dimensional chain with a quasiperiodic on-site potential, mimicking a quasicrystal compound. That is, the potential is not random, as in the Anderson model, but it is not periodic, as required for extended Bloch states Aulbach et al. (2004); Bu et al. (2022). As discussed below, the AAH model is notable for being self-dual, thus having extended-localized phase transitions to all states at the same critical point Domínguez-Castro and Paredes (2019). Therefore, the main goal of this paper is to investigate how the cavity QED affects the critical properties of the AAH model, in particular in the strong coupling regime.
@@
We analyze the cavity QED Aubry-Andre-Harper model by exact diagonalization methods, examining localization properties by the behavior of the inverse participation ratio and multifractal analysis. We show that the critical point that exists in the absence of the cavity changes into an entire critical phase where critical, extended, and localized eigenstates coexist, and that broadens with increasing cavity coupling.
@@
This paper is organized as follow. The AAH model, its coupling with the electromagnetic field of the cavity and the quantities of interest are presented in Section II. In Section III, we present and discuss our main results, while our conclusions and further remarks are left to Section IV.
@@
Figure 1: (a) Schematic representation of a chain inside an optical cavity, where the heights indicate the number of photons, and $\gamma$ represents the coupling with the electromagnetic field. (b) As the electron hops, it can absorb or emit a photon, effectively leading to a multi-chain perspective, with |$\ket{n}$ denoting the photon state.
@@ II.1 The model
$+g_{AA}\sum_{\mathbf{i}}\cos{(2\pi\beta i+\phi)}c^{\dagger}_{\mathbf{i}}c_{\mathbf{i}}^{\phantom{\dagger}},$ (1)
where the sums run over a one-dimensional chain, with $\langle\mathbf{i},\mathbf{j}\rangle$ denoting nearest neighbor sites. Here, we use the standard second quantization formalism, in which $c_{\mathbf{i}}^{\dagger}$ ($c_{\mathbf{i}}$) are creation (annihilation) operators of fermions at a given site $\mathbf{i}$. The first term on the right-hand side of Eq. (1) describes the fermionic hopping, while the second term corresponds to the onsite potential, with strenght $g_{AA}$. To describe incommensurate features, one may deal with open boundary conditions (OPC) and define $\beta=\frac{\sqrt{5}-1}{2}$, i.e. as the inverse golden ratio, with $\phi$ being an arbitrary global phase. For periodic boundary conditions (PBC), it is achieved by defining $\beta$ as the ratio of two adjacent Fibonacci numbers, $F_{m-1}/F_{m}$, and extrapolating the results to the thermodynamic limit. Hereafter, the hopping integral, $t$, sets the energy scale and may be taken as unity; similarly, we also take the lattice constant, $a$, as unity.
@@
As mentioned in the Introduction, we place our one-dimensional AAH electronic system within a cavity with a single-mode electromagnetic field; see Fig. 1. Following Refs. Svintsov et al. (2024) and Li et al. (2020), the coupling between the bosonic and fermionic degrees of freedom changes the hopping integral (Peierls substitution) Peierls (1933), $t\to t\exp[ieaA_{x}/\hbar c\big{]}$, with $A_{x}$ being the vector potential of the electromagnetic field, and $e$ the fermion charge. Here we assume that the wavelength is much larger than the lattice spacing, leading to a position-independent vector potential. Given this, we define $A_{x}=A_{0}(\hat{b}+\hat{b}^{\dagger})$, with $\hat{b}^{\dagger}$ ($\hat{b}$) being creation (annihilation) operators of photons of frequency $\omega_{0}$. The cavity-coupled Hamiltonian then reads,
@@
$+g_{AA}\sum_{\mathbf{i}}\cos{(2\pi\beta i+\phi)}c^{\dagger}_{\mathbf{i}}c_{\mathbf{i}}^{\phantom{\dagger}}+\hbar\omega_{0}\hat{b}^{\dagger}\hat{b},$ (2)
@@
We investigate the Hamiltonian of Eq. (2) through exact diagonalization methods in the subspace of a single fermion coupled to $N_{ph}$ photons. The operators $e^{i\gamma(\hat{b}+\hat{b}^{\dagger})}$ are obtained analytically through the Baker-Campbell-Hausdorff formula (see, e.g., Ref. Svintsov et al. (2024)). In this work, we use both OBC and PBC, while averaging over 10 different random values of $\phi$. We also set a cutoff for the total number of photons at $N_{ph}=8$, which is sufficiently large to ensure that the inclusion of more photon will correspond to weak (negligible) corrections in the inverse participation ratio.
@@ II.2 The inverse participation ratio
A convenient basis for our Hilbert space is spanned by |$\ket{l,n}$, with $1\leq l\leq L$ labeling the site position, and $0\leq n\leq N_{ph}$ being the number of photons. We may therefore write a generic (normalized) eigenstate (normalized) of $\mathcal{H}$ as
@@
where $j$ labels the eigenstate, and $\phi_{l,n}^{(j)}$ is the probability amplitude of finding the system in state |$\ket{l,n}$.
Given this, one key quantity of interest is the inverse participation ratio (IPR), which determines if a given wavefunction is spatially localized; it is defined as
@@
with $q=2$. If $\text{IPR}(\ket{\psi_{j}})\to 1$ when $L\to\infty$, the state |$\ket{\psi_{j}}$ is localized at a single site. Otherwise $\text{IPR}(\ket{\psi_{j}})\propto 1/N$ so that |$\ket{\psi_{j}}$ extends over $N$ sites.
@@
Similarly, and within the same energy window, one may define the minimum IPR, that is
@@
While $\overline{\text{IPR}}$ determines the occurrence of localized states for a given set of external parameters, $\text{IPR}_{min}$ establishes the threshold for the occurrence of at least one extended state. Unless otherwise mentioned, in this work we examine both IPR’s for the entire energy spectrum.
Figure 2: Numerical energy spectrum $E$ as a function of $g_{AA}$ for $L=300$ sites and fixed $\phi=0$. Panel (a) shows the standard AAH model (i.e. no cavity effects), while panel (b) presents the case with coupling with photons. For the latter, we fixed $\omega_{0}=1$ and $\gamma=0.15$, for $N_{ph}=8$ photons. The heatmap describes the IPR associated to each eigenstate, plotted in 10-base logarithmic scale.
II.3 Fractal properties
@@
Another interesting feature of the IPR is its relationship with the fractal dimension of the wavefunction, which indicates how distributed it is within the medium. For extended states, the probability of finding the particle at a given site is proportional to $1/L^{d}$, where $d$ is the dimensionality of the system. This leads to $\text{IPR}\approx L^{d(1-q)}$, which simplifies to $\text{IPR}\approx L^{-d}$ for $q=2$. However, at the critical point, the wavefunction coverage over the medium becomes fractal, rather than homogeneous; accordingly, the scaling properties of the IPR are described through the replacement $d\to D_{f}$, where $D_{f}$ is the fractal dimension.
In the presence of $N_{ph}$ photons, the extension of these ideas is straightforward. Since one expects that $|\phi_{l,n}^{j}|^{2}\approx 1/(L\times N_{ph})^{d}$ for extended states, then $\text{IPR}\approx(L\times N_{ph})^{d(1-q)}$, so that
$\log(\text{IPR})\approx-D_{f}\big{[}\log(L)+\log(N_{ph})\big{]},$ (7)
for $q=2$. Therefore, for large, but finite $N_{ph}$, $D_{f}$ may still be extracted from a double-logarithm analysis of the IPR.
@@
Further features of the wave function are provided by performing a multifractal analysis Kohmoto (1983); Liu et al. (2022). Assuming PBC, we choose system sizes from a Fibonacci sequence, $L=F_{m}$. As the probability of finding the particle on a given site $l$ is $P_{l}\propto F_{m}^{-\alpha_{l}^{(j)}}$, the set of exponents $\alpha_{l}^{(j)}$ are used to characterize the distribution of the state |$\ket{\psi_{j}}$ over the sites, as follows. We define $\alpha_{min}^{(j)}\equiv\text{min}\big{[}\alpha_{l}^{(j)}\big{]}$, with $1\leq l\leq F_{m}$, and take $m\to\infty$: if $\alpha_{min}^{(j)}\to 1$ or 0, the wave function is extended or localized, respectively; if $10<\alpha_{min}^{(j)}<$ 1, it is fractal.
@@
Similarly to the IPR case, it is useful to perform a multifractal analysis. With the same definitions as those leading to Eq. (5), we consider the averages
@@
Figure 3: Inverse participation ratio results of the standard AAH model. (a) $\overline{\text{IPR}}$ as a function of $g_{AA}$ for different system sizes $L$. (b) $\overline{\text{IPR}}$ as a function of $1/L$ (symbols) and extrapolated to the thermodynamic limit (solid lines). (c) IPRmin as a function of $g_{AA}$ for different system sizes $L$. (d) IPRmin as a function of $1/L$ (symbols) and extrapolated to the thermodynamic limit (solid lines). Here, and in all subsequent figures, when not shown, error bars are smaller than symbol size.

### v2

@@ Abstract
We theoretically investigate criticality and multifractal states in a one-dimensional Aubry-Andre-Harper model coupled to electromagnetic cavities. We focus on two specific cases where the phonon frequencies are $\omega_{0}=1$ and $\omega_{0}=2$, respectively. Phase transitions are analyzed using both the average and minimum inverse participation ratio to identify metallic, fractal, and insulating states. We provide numerical evidence to show that the presence of the optical cavity induces a critical, intermediate phase in between the extended and localized phases, hence drastically modifying the traditional transport phase diagram of the Aubry-Andre-Harper model, in which critical states can only exist at the well-defined metal-insulator critical point. We also investigate the probability distribution of the inverse participation ratio and conduct a multifractal analysis to characterize the nature of the critical phase, in which we show that extended, localized, and fractal eigenstates coexist. Altogether our findings reveal the pivotal role that the coupling to electromagnetic cavities plays in tailoring critical transport phenomena at the microscopic level of the eigenstates.
@@ I Introduction
In view of these stimulating results, a great theoretical effort has been made over the past years, to understand the leading effects of non-perturbative cavity light-matter coupling. For instance, Ref. Hagenmüller et al. (2017) showed that the coupling with the cavity enhances the conductivity. Further, a Dicke superradiant quantum phase transition may emerge in a cavity coupled with a 2D electron gas with Rashba spin-orbit interaction Nataf et al. (2019). Other important theoretical developments were also made for topological phases Wang et al. (2019); Guerci et al. (2020); Jangjan and Hosseini (2020); Dmytruk and Schirò (2022); Allard and Weick (2023); Ezawa (2024); Liu et al. (2023); Nguyen et al. (2024), Majorana fermions Bacciconi et al. (2024); Gómez-León et al. (2024), entanglement properties Passetti et al. (2023), magnetic or superconducting properties Ashida et al. (2020); Masuki and Ashida (2024); Li and Eckstein (2020), Kondo effect Mochida and Ashida (2024), and disordered systems Arwas and Ciuti (2023); Svintsov et al. (2024); Guo and Cai (2024). In particular, Ref. Svintsov et al. (2024) investigated the effects of cavity QED in an insulating 1D disordered chain, demonstrating that conductivity can be enhanced by several orders of magnitude due to the emission and absorption of virtual photons. This implies that the localization length is highly dependent on the coupling with the cavity Svintsov et al. (2024). Since the 1D Anderson model does not exhibit a metal-insulator transition, these results only apply to exponentially localized, Anderson wavefunctions in disordered media. However, the impact of the coupling to electromagnetic cavities on metal-insulator transitions remains unknown.
@@
To address this issue, we consider a simple Hamiltonian exhibiting extended-localized phase transition, the Aubry-Andre-Harper (AAH) model Harper (1955); Aubry and André (1980). It describes a one-dimensional chain with a quasiperiodic on-site potential, mimicking a quasicrystal. That is, the potential is not random, as in the Anderson model, but aperiodic, as required for extended Bloch states Aulbach et al. (2004); Bu et al. (2022). As discussed below, the AAH model is notable for being self-dual, thus having extended-localized phase transitions to all states at the same critical point Domínguez-Castro and Paredes (2019). Therefore, the main goal of this paper is to investigate how the cavity QED affects the critical properties of the AAH model, particularly in the strong coupling regime.
@@
We analyze the cavity QED Aubry-Andre-Harper model through exact diagonalization methods; localization properties are probed by the behavior of inverse participation ratio and multifractal analysis. We show that the critical point that exists in the absence of the cavity changes into an entire critical phase where critical, extended, and localized eigenstates coexist, which broadens with increasing cavity coupling.
@@
The paper is organized as follows. The AAH model, its coupling with the electromagnetic field of the cavity and the quantities of interest are presented in Section II. In Section III we present and discuss our main results, while our conclusions and further remarks are left to Section IV.
@@
Figure 1: (a) Schematic representation of a chain inside an optical cavity. The horizontal lines lie on the chain sites, and denote photon levels, whose height indicates the number, $n$, of electromagnetic quanta. $\gamma$ is the coupling strength with the electromagnetic field within the cavity. (b) As the electron hops, it can absorb or emit a photon, effectively leading to a multi-chain perspective, with |$\ket{n}$ denoting the photon state.
@@ II.1 The model
$+g_{AA}\sum_{\mathbf{i}}\cos{(2\pi\beta|\mathbf{i}|+\phi)}c^{\dagger}_{\mathbf{i}}c_{\mathbf{i}}^{\phantom{\dagger}},$ (1)
where the sums run over a one-dimensional chain. Here, we use the standard second quantization formalism, in which $c_{\mathbf{i}}^{\dagger}$ ($c_{\mathbf{i}}$) are creation (annihilation) operators of fermions at a given site $\mathbf{i}$. The first term on the right-hand side of Eq. (1) describes the fermionic hopping, while the second term corresponds to the onsite potential, with strenght $g_{AA}$. To describe incommensurate features, one may deal with open boundary conditions (OPC) and define $\beta=\frac{\sqrt{5}-1}{2}$, i.e. as the inverse golden ratio, with $\phi$ being an arbitrary global phase. For periodic boundary conditions (PBC), it is achieved by defining $\beta$ as the ratio of two adjacent Fibonacci numbers, $F_{m-1}/F_{m}$, and extrapolating the results to the thermodynamic limit. Hereafter, the hopping integral, $t$, sets the energy scale and may be taken as unity; similarly, we also take the lattice constant, $a$, as unity.
@@
As mentioned in the Introduction, we place our one-dimensional AAH electronic system within a cavity with a single-mode electromagnetic field. Following Refs. Svintsov et al. (2024) and Li et al. (2020), the coupling between the bosonic and fermionic degrees of freedom changes the hopping integral (as a Peierls substitution) Peierls (1933); Marder (2010), $t\to t\exp[ieaA_{x}/\hbar c\big{]}$, with $A_{x}$ being the vector potential of the electromagnetic field, and $e$ the fermion charge. Here we assume that the wavelength is much larger than the lattice spacing, leading to a position-independent vector potential. Given this, we define $A_{x}=A_{0}(\hat{b}+\hat{b}^{\dagger})$, with $\hat{b}^{\dagger}$ ($\hat{b}$) being creation (annihilation) operators of photons of frequency $\omega_{0}$. This is illustrated in Fig. 1 (a), where the renormalized (Peierls) electronic hopping connects states with different photon numbers through absorption or emission processes. The cavity-coupled Hamiltonian then reads,
@@
$+g_{AA}\sum_{\mathbf{i}}\cos{(2\pi\beta|\mathbf{i}|+\phi)}c^{\dagger}_{\mathbf{i}}c_{\mathbf{i}}^{\phantom{\dagger}}+\hbar\omega_{0}\hat{b}^{\dagger}\hat{b},$ (2)
@@
We investigate the Hamiltonian of Eq. (2) through exact diagonalization methods in the subspace of a single fermion coupled to $N_{ph}$ photons. A convenient basis for our Hilbert space is spanned by |$\ket{l,n}$, with $1\leq l\leq L$ labeling the site position, and $0\leq n\leq N_{ph}$ being the photon quantum number. Using the Baker-Campbell-Hausdorff formula, the matrix elements for the Peierls operator can be expressed as
$\langle n|e^{i\gamma(b+b^{\dagger})}|m\rangle=$ $e^{-\gamma^{2}/2}\sum_{s=0}^{n}\sum_{j=0}^{m}\frac{(i\gamma)^{s}}{s!}\frac{(i\gamma)^{j}}{j!}\times\delta_{n-s,m-j}P(s,n)P(j,m)~,$ (3)
obtained by expanding the operators as a Taylor series, and with $P(j,m)=\sqrt{m[m-1][m-2]\dots[(m-(j-1)]}$ (see, e.g., Ref. Svintsov et al. (2024)). As mentioned before, it renormalizes the hopping integrals within the different photon processes. Given this, the matrix representation of the Hamiltonian is
$\mathcal{H}=\begin{pmatrix}\hat{H}_{0,0}&\hat{H}_{0,1}&\dots&\hat{H}_{0,N_{ph}}\\ \hat{H}_{1,0}&\hat{H}_{1,1}&\dots&\hat{H}_{1,N_{ph}}\\ \vdots&\vdots&\dots&\vdots\\ \hat{H}_{N_{ph},0}&\hat{H}_{N_{ph},1}&\dots&\hat{H}_{N_{ph},N_{ph}}\end{pmatrix}$
with $\hat{H}_{n,m}$ denoting $L\times L$ matrices whose matrix elements are $\big{[}\hat{H}_{n,m}\big{]}_{i,j}=\langle i,n|\mathcal{H}|j,m\rangle$. The final $(N_{ph}+1)\times L$ matrix is fully diagonalized by standard linear algebra numerical routines, which provide the eigenvalues and eigenvectors.
This scenario effectively corresponds to a problem involving strongly coupled chains, as illustrated in Fig. 1 (b). In this representation, |$\ket{n}$ denotes the photon number state, and the links between different photon states correspond to photon-assisted hopping terms. In words, the original system without photons, described by $\hat{H}_{0,0}$, is coupled to a series of virtual replicas – each with $n$ photons, renormalized hopping amplitudes, and energy spectra shifted by $n\hbar\omega_{0}$. These replicas are represented by the matrices $\hat{H}_{n,n}$. The interaction between these chains arises through photon absorption and emission processes, which are captured by the off-diagonal matrices $\hat{H}_{n,m}$ for $n\neq m$. In this work, we use both OBC and PBC, while averaging over 10 different random values of $\phi$. The relatively small number of disorder configurations is justified by the weak dependence of the inverse participation ratio on $\phi$ in the AAH model, as confirmed a posteriori. We also set a cutoff for the total number of photons at $N_{ph}=8$, as discussed below.
@@ II.2 The inverse participation ratio
Using the basis states |$\ket{l,n}$ for our Hilbert space, we may write a generic (normalized) eigenstate of $\mathcal{H}$ as
@@
where $j$ labels the number of the eigenstate, and $\phi_{l,n}^{(j)}$ is the probability amplitude of finding the system in |$\ket{l,n}$.
Given this, one key quantity of interest is the inverse participation ratio (IPR), which determines if a given wavefunction is spatially localized. It is defined as
@@
with $q=2$. If $\text{IPR}(\ket{\psi_{j}})\to 1$ in the limit $L\to\infty$, the state |$\ket{\psi_{j}}$ is localized in a single site. On the other hand, if $\text{IPR}(\ket{\psi_{j}})\to 0$ in the limit $L\to\infty$, the state is extended, meaning that |$\ket{\psi_{j}}$ is delocalized and spreads over all $L$ sites in the system.
Figure 2: Log-linear plots of the relative $\overline{\text{IPR}}$ error [Eq. (10)], averaged over the energy ranges (a) $\Delta E=[1.5,2.5]$, and (b) $\Delta E=[3.5,4.5]$, as functions of the number of photons, $N_{\text{ph}}$, for different values of the disorder strength, $g_{AA}$.
@@
We note that the emission and absorption of photons couple different states across the photon-number subspaces, which would lead to inhomogeneities for the localization properties within a given energy window. Therefore, it is useful to define the typical IPR,
$\text{IPR}_{\text{typ}}=e^{\langle\ln\text{IPR}\rangle}~,$ (7)
with
$\langle\ln\text{IPR}\rangle=\frac{1}{N_{st}}\sum_{j=M_{i}}^{M_{f}}\ln[\text{IPR}(\ket{\psi_{j}})\big{]}~.$ (8)
This quantity effectively distinguishes delocalized from localized states and provides a more accurate determination of the critical points. Similarly, and within the same energy window, one may define the minimum IPR, that is
@@
While $\overline{\text{IPR}}$ and $\text{IPR}_{\text{typ}}$ determine the occurrence of localized states for a given set of external parameters, $\text{IPR}_{min}$ establishes the threshold for the occurrence of at least one extended state.
II.3 The cutoff in the number of photons $N_{ph}$
At this point, it is important to examine the cutoff of $N_{\text{ph}}$, which should be chosen such that it does not affect the accuracy of the physical quantities of the system. Since the different photon-number subspaces $\hat{H}_{n,n}$ are energetically shifted by $n\hbar\omega_{0}$, the impact of the cutoff depends on the specific energy window being investigated. Of particular interest are the low-energy states, due to their importance for transport properties. One might expect these states to be only weakly affected by contributions from subspaces with large number of photons, especially when $n\hbar\omega_{0}$ exceeds the bare electron bandwidth.
To determine the cutoff for the photon number $N_{\text{ph}}$, we analyze the behavior of $\overline{\text{IPR}}$ as a function of $N_{\text{ph}}$. This is done by comparing it to a reference case with a sufficiently large photon number, which serves as an approximation for the asymptotic behavior. To quantify the deviation from the reference, we compute the relative error in $\overline{\text{IPR}}$,
$\delta(N_{\text{ph}})=\frac{|\overline{\text{IPR}}(N_{\text{ph}})-\overline{\text{IPR}}_{\text{ref}}|}{\overline{\text{IPR}}_{\text{ref}}}~,$ (10)
with $\overline{\text{IPR}}_{\text{ref}}$ being the reference value taken from the case with $N_{\text{ph}}=12$. The behavior of $\delta(N_{\text{ph}})$ is shown in Fig. 2 for two different energy windows. In both scenarios, and across different disorder strengths, the corrections to $\overline{\text{IPR}}$ become negligible once the photon number reaches $N_{\text{ph}}\geq 8$. Although higher-energy states might require more photons for an accurate representation, such states are generally less significant in the context of transport experiments. Therefore, in what follows, we fix the photon number to $N_{\text{ph}}=8$.
II.4 Fractal properties
@@
Another interesting feature of the IPR is its relation with the fractal dimension of the wavefunction, which indicates how distributed it is within the medium. For extended states, the probability of finding the particle at a given site is proportional to $1/L^{d}$, where $d$ is the dimensionality of the system. This leads to $\text{IPR}\approx L^{d(1-q)}$, which simplifies to $\text{IPR}\approx L^{-d}$ for $q=2$. However, at the critical point, the wavefunction coverage over the medium becomes fractal, rather than homogeneous; accordingly, the scaling properties of the IPR are described through the replacement $d\to D_{f}$, where $D_{f}$ is the fractal dimension.
In the presence of $N_{ph}$ photons, the extension of these ideas is straightforward. For extended states, the wavefunction amplitudes are expected to scale as $|\phi_{l,n}^{j}|^{2}\approx A_{n}/L^{d}$, where $A_{n}$ is a normalization factor that depends on the photon subspace index $n$. Under this assumption, the IPR behaves as
$\text{IPR}\approx\left(\sum_{n=1}^{N_{ph}}A_{n}^{q}\right)L^{d(1-q)}=C_{N_{ph}}L^{d(1-q)}~,$
with $C_{N_{ph}}=\sum_{n=1}^{N_{ph}}A_{n}^{q}$ taking into account the contributions from different photon subspaces. This leads to the scaling relation
$\log(\text{IPR})\approx-D_{f}\log(L)+\log(C_{N_{ph}}),$ (11)
for $q=2$. Importantly, for large photon number cutoffs $N_{ph}$, the sum over $n$ in $C_{N_{ph}}$ is often dominated by a few photon subspaces, rendering $C_{N_{ph}}$ constant. As a result, $C_{N_{ph}}$ does not alter the scaling behavior of the IPR with system size, thus not affecting the extracted fractal dimension $D_{f}$.
@@
Further features of the wavefunction are provided by performing a multifractal analysis Kohmoto (1983); Liu et al. (2022). Assuming PBC, we choose system sizes from a Fibonacci sequence, $L=F_{m}$. As the probability of finding the particle on a given site $l$ is $P_{l}\propto F_{m}^{-\alpha_{l}^{(j)}}$, the set of exponents $\alpha_{l}^{(j)}$ are used to characterize the distribution of the state |$\ket{\psi_{j}}$ over the sites, as follows. We define $\alpha_{min}^{(j)}\equiv\text{min}\big{[}\alpha_{l}^{(j)}\big{]}$, with $1\leq l\leq F_{m}$, and take $m\to\infty$: if $\alpha_{min}^{(j)}\to 1$ or 0, the wave function is extended or localized, respectively; if $0<\alpha_{min}^{(j)}<1$, it is fractal.
@@
Similarly to the IPR case, it is useful to perform a multifractal analysis within a given energy window. Then, with the same definitions as those leading to Eq. (6), we consider the averages
@@
Figure 3: Numerical energy spectrum $E$ as a function of $g_{AA}$ for $L=300$ sites and fixed $\phi=0$. Panel (a) shows the standard AAH model (i.e. no cavity effects), while panel (b) presents the case with coupling with photons. For the latter, we fixed $\omega_{0}=1$ and $\gamma=0.15$, for $N_{ph}=8$ photons. The heatmap describes the IPR associated with each eigenstate, plotted in 10-base logarithmic scale.